One positive number is 5 times another number. The difference between the two numbers is 124, find the numbers.
Found 3 solutions by scott8148, stanbon, bucky:
Answer by scott8148(6628)

let x=number, so 5x="other number" ... 5x-x=124

Answer by stanbon(75887)

One positive number is 5 times another number. The difference between the two numbers is 124, find the numbers.
------------------------------------------
EQUATIONS:
x = 5y
x-y =124
---------------------
Substitute into x-y=124 to solver for y:
5y-y=124
4y = 124
y = 31
-------------
Substite into x=5y to solve for "x":
x = 5*31 = 155
====================
Cheers,
Stan H.

Answer by bucky(2189)

Let one of the numbers be x. You are told that the other number is 5 times x or 5x. The
difference between the numbers is 5x - x which equals 4x. But you are also told that this
difference between the two numbers is 124. So we have two values for the difference in
the numbers and in equation form this can be written as:
.
4x = 124
.
Solve for x by dividing both sides by 4 (which is the multiplier of the x) to get:
.
x = 124/4 = 31
.
You now know that the smaller of the two numbers is 31. Five times this smaller number gives
you the larger number. Therefore, the larger number is 5 * 31 = 155.
.
You can check these two numbers by subtracting them and finding that the difference is:
.
155 - 31 = 124
.
This difference in the numbers agrees with the value provided in the problem ... so our answer
is correct ... the two numbers are 31 and 155.
.
Hope this helps you to understand the problem and how to work toward getting the answer.
.